OpenStudy (anonymous):

how do i simplify e^ln3

OpenStudy (anonymous):

The e and ln cancel, leaving just 3

OpenStudy (anonymous):

ne^1

OpenStudy (anonymous):

how do i simplify e^(1+lnx)

OpenStudy (anonymous):

actually.... e^(1+lnx) = e^1*e^lnx = e *x = ex

OpenStudy (anonymous):

That's better

OpenStudy (anonymous):

Bliss can I have help on my latest question?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is there any way i can simplify square root of x/x any more

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

\[\frac{\sqrt{x}}{x}\] Is this what you're trying to simplify?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You cannot simpify it anymore....Are you sure that's the entire problem?

OpenStudy (anonymous):

It's not set equal to anything?

OpenStudy (anonymous):

Bliss may I please have help again?

OpenStudy (anonymous):

no thats the entire problem

OpenStudy (anonymous):

You do not need to simplify that any further that I know of. If the root was in the denominator, it would be a different story